# High-dimensional sliced inverse regression with endogeneity

### v1

@@ 4.1 First stage results
In the first stage of the 2SLSIR estimation procedure, each column of the matrix $\bm{\Gamma}$ is a Lasso estimator from the regression of $\mathbf{x}_{j}$ on the design matrix $\mathcal{Z}$. We obtain the following result as a straightforward application of Banerjee et al., (2014); all proofs are provided in the Supplementary Materials S2.
@@
Assume Conditions 4.1 and 4.2 are satisfied, and set
@@
Then with probability at least $1-\exp(-C^{\prime}\log q)$, the columns of $\widehat{\bm{\Gamma}}$ satisfy
$\left\lVert\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\right\rVert_{1}\leq C\sigma_{\max}r\{\log(pq)/n\}^{1/2}.$ It follows that with probability at least $1-\exp(-C^{\prime}\log pq)$, we obtain
@@ 4.2 Second stage results
In the second stage, we compute the Lasso SIR estimator using the outcome $y$ and the values $\widehat{\mathcal{X}}=\mathcal{Z}\widehat{\bm{\Gamma}}$ obtained from the first stage. We require additional technical assumptions.
Condition 4.4.
@@
There exists a constant $c$ such that $\lambda_{d}\geq c>0$, where $\lambda_{d}$ is the $d$th largest eigenvalue of $\bm{\Lambda}=\operatorname{\text{Cov}}\left\{E\left(\widehat{\mathbf{X}}\mid y\right)\right\}$.
Condition 4.5.
The covariance matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$ satisfies the RE condition with parameters ($\kappa_{1},3)$, for some constant $\kappa_{1}>0$ and $\|\bm{\Sigma}_{\widehat{\mathbf{X}}}\|_{\max}\leq C<\infty$
Condition 4.6.
Let $y_{(1)},\ldots,y_{(n)}$ be the order statistics of the sample response data $y_{1},\ldots,y_{n}$. Let $C>0$ and let $\Pi_{n}(C)$ denote the collection of all the $n$-point partitions $-C\leq y_{(1)}\leq\cdots\leq y_{(n)}\leq C$ on the interval $[-C,C]$. We assume the vector-valued random variable $\bm{m}(y_{(i)})=E(\mathbf{Z}\mid y_{(i)})$ to have a total variation of order $1/4$. That is, for any fixed $C>0$, we have
$\lim_{n\rightarrow\infty}\frac{1}{n^{1/4}}\sup_{\Pi_{n}(C)}\sum_{i=2}^{n}\left\|\bm{m}\left\{y_{(i)}\right\}-\bm{m}\left\{y_{(i-1)}\right\}\right\|_{\infty}=0.$
@@
Condition 4.4 is analogous to the coverage condition imposed in the SIR literature for $\bm{\Lambda}$ (e.g., Li,, 1991). Condition 4.5 ensures the loss function corresponding to the second stage of the proposed Lasso SIR estimator does not have too small curvature, and is similar to a condition imposed by Lin et al., (2015) for the two-stage penalized least square method. Finally, Condition 4.6 controls the smoothness of the conditional expectation $\bm{m}(y)=E(\mathbf{Z}\mid y)$. Such a similar condition is used in Tan et al., (2018) when no endogeneity is present, and similar conditions on the total variation have been introduced in much earlier work on SIR (e.g., Hsing and Carroll,, 1992; Zhu et al.,, 2006).
@@
We first consider the case of $d=1$ i.e., the single index model version of (2). Here, Lemma 1 implies the covariance matrix $\bm{\Lambda}=\operatorname{\text{Cov}}\left\{E\left(\widehat{\mathbf{X}}\mid y\right)\right\}$ has rank $d=1$ and hence can be written in the form $\bm{\Lambda}=\lambda\bm{\eta}\bm{\eta}^{\top}$, where $\lambda>0$ denotes the non-zero eigenvalue of $\bm{\Lambda}$ and $\bm{\eta}$ is the corresponding eigenvector. Without loss of generality, assume $\|\bm{\beta}\|=1$. For convenience below, we omit the subscript $k$ and use $\hat{\lambda}$ and $\hat{\bm{\eta}}$ to denote the largest eigenvalue and its corresponding eigenvector of $\hat{\bm{\Lambda}}$, respectively, noting both were defined above (4). We also write $\tilde{\mathbf{y}}_{1}=\tilde{\mathbf{y}}$. With this notation, the 2SLSIR estimator for the column space of $\bm{\beta}$ is given by
@@
By Lemma (1), we have $\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}=\xi\bm{\eta}$, where $\xi=\left\lVert\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}\right\rVert_{2}>0$ is a constant. Letting $\tilde{\bm{\eta}}=P(\bm{\eta})\hat{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\hat{\bm{\eta}}=\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}$, where $\tilde{\bm{\beta}}=\xi^{-1}\bm{\beta}{\bm{\eta}}^{\top}\hat{\bm{\eta}}$, then it is straightforward to see $\tilde{\bm{\beta}}$ has the same direction as $\bm{\beta}$, i.e., $\mathcal{P}(\tilde{\bm{\beta}})=\mathcal{P}(\bm{\beta})$, when $\bm{\eta}^{\top}\hat{\bm{\eta}}\neq 0$. The following result guarantees $\bm{\eta}^{\top}\hat{\bm{\eta}}\neq 0$ holds with high probability.
@@
Let $\hat{\rho}=\bm{\eta}^{\top}\hat{\bm{\eta}}$, and assume Conditions 4.1-4.6 are satisfied. Then with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$, it holds that
$\hat{\rho}^{2}\geq 1-\frac{CL}{\lambda}\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}.$
We now proceed to condition on ${\bm{\eta}}^{\top}\hat{\bm{\eta}}\neq 0$ and an RE property for the sample covariance $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}$.
Lemma 3.
Suppose we condition on $\hat{\rho}={\bm{\eta}}^{\top}\hat{\bm{\eta}}\neq 0$ and the sample covariance matrix $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}$ satisfies an RE property with parameters $(\kappa_{2},3)$; see Proposition S2 in Supplementary Material S2 which shows the latter holds with high probability. If the tuning parameter satisfies $\mu_{2}\geq 2\left\|\hat{\bm{\eta}}-n^{-1}\hat{\mathcal{X}}^{\top}\hat{\mathcal{X}}\tilde{\bm{\beta}}\right\|_{\infty}$, then any solution $\hat{\bm{\beta}}$ defined in (8) satisfies
$\left\lVert\hat{\bm{\beta}}-\tilde{\bm{\beta}}\right\rVert_{2}\leq\dfrac{12}{\kappa_{2}}\mu_{2}\sqrt{s},$
We emphasize that the theoretical analyses for our proposed 2SLSIR estimator do not require the covariance matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}$ to have full rank, meaning the results in this section still hold even when $q<p$. Following Lemma 3, we obtain the following main result on the consistency of the 2SLSIR estimator for the single index model.
@@
Assume Conditions 4.1-4.6 are satisfied, and choose the first stage tuning parameters $\mu_{1j}$ as in (6). If we further select the second stage tuning parameters to satisfy
$\mu_{2}=C\left(\sqrt{\dfrac{L^{*}}{\lambda}}+L^{*}\right)\;\text{with}\;L^{*}=L\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$ (9)
then with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$, the 2SLSIR estimator $\hat{\bm{\beta}}$ satisfies
$\left\|P(\hat{\bm{\beta}})-P({\bm{\beta}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\sqrt{\dfrac{L^{*}}{\lambda}}+L^{*}\right).$
Compared with the bound for the one-stage Lasso SIR estimator established by Lin et al., (2019), the theoretical bound for the 2SLSIR estimator has an additional factor of $\sigma_{\max}r\left\{\log(pq)/n\right\}^{1/2}$, reflecting the consequences of endogeneity. Moreover, compared with the result for the two-stage penalized least squares estimator in the linear outcome model of Lin et al., (2015), the bound in Lemma 2 involves an additional term depending upon the largest eigenvalue $\lambda$ of $\operatorname{\text{Cov}}\left\{E(\widehat{\mathbf{X}}\mid y\right\}$, reflecting the consequences of non-linearity. Intuitively, the estimation error of $\hat{\bm{\beta}}$ depends on the estimation error of $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$, which in turn depends upon the estimation error of both the coefficients $\hat{\bm{\Gamma}}$ in the first stage and the covariance matrix $\bm{\Sigma}_{\mathbf{Z}}$. Both $p$ and $q$ enter this bound logarithmically.
Next, we establish the selection consistency of the 2SLSIR estimator. Let $\hat{S}=\{j:\hat{\beta}_{j}\neq 0\}$, and partition the matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$ as
@@
Let $\phi=\|(\mathbf{C}_{SS})^{-1}\|_{\infty}$, and note in this case we have $\xi=\left\lVert\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}\right\rVert_{2}=\left\lVert\mathbf{C}_{SS}\bm{\beta}_{S}\right\rVert_{2}$.
Condition 4.7.
@@
Condition 4.7 is similar to the mutual incoherence condition for the ordinary Lasso estimator in the linear model without endogeneity (e.g., see Section 7.5.1, Wainwright,, 2019) and is standard in many high-dimensional regression contexts Li et al., (2017).
@@
Assume Conditions 4.1-4.7 are satisfied, and we choose the first stage tuning parameters $\mu_{1j}$ as in (6). If there exists a constant $C$ such that
@@
and the second stage tuning parameters are also selected to satisfy
$\min_{j\in S}|\beta_{j}|>\frac{2\xi\hat{\rho}}{2-\alpha}\varphi\mu_{2},$ (11)
then with probability at least $1-\exp(C^{\prime}\log q)-\exp(C^{\prime\prime}\log pq)$, there exists a solution $\hat{\bm{\beta}}$ of (8) whose support satisfies $\hat{S}=S$.
Since the mutual incoherence condition 4.7 is imposed for the population quantity $\bm{\Sigma}_{\widehat{\mathbf{X}}}$, the additional assumption in (10) is necessary to ensure that this condition is satisfied for the sample covariance matrix $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}$ with high probability. Furthermore, (11) ensures that the minimum non-zero value $\tilde{\beta}_{j}$ can be identified, and has an additional factor $\xi\hat{\rho}$ in the numerator, reflecting another consequence of non-linearity. Note the lower bound is small when $\alpha$ is small, that is, the lower the correlation between the important and non-important variables, the smaller the magnitude of non-zero components which can be identified.
@@ 4.3 Extension to multiple index models
Finally, we generalize the above results to multiple index models. For $k=1,\ldots,d$, let $\widehat{\mathcal{S}}_{k}=\{j:|\hat{\beta}_{jk}|\neq 0\}$ be the estimated support set for the vector $\hat{\bm{\beta}}_{k}$, and $\widehat{\mathcal{S}}=\cup_{k}{\mathcal{S}_{k}}$. Recall that $\mathbf{B}$ is only identifiable up to rotational transformation, and the latter only preserves the zero rows of $\mathbf{B}$, then selection in the multiple index models refers only to selecting non-zero rows of $\mathbf{B}$. Equivalently, we wish that the set $\hat{\mathcal{S}}$ consistently estimates the support set $\mathcal{S}$ of the diagonal of $\mathcal{P}(\mathbf{B})$.
@@
Assume Conditions 4.1-4.6 are satisfied, and choose the first stage tuning parameters $\mu_{1j}$ as in (6). If the second stage tuning parameters $\mu_{2k}$ are also selected to satisfy
$\mu_{2k}=C\left(\sqrt{\dfrac{L^{*}}{\lambda_{d}}}+L^{*}\right)\;\text{where}\;L^{*}=L\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$
for $k=1,\ldots,d$:
@@
    Then with probability at least $1-d\exp(-C^{\prime}\log q)-d\exp(-C^{\prime\prime}\log pq)$, the 2SLSIR estimator $\widehat{\mathbf{B}}=(\hat{\bm{\beta}}_{1},\ldots,\hat{\bm{\beta}}_{d})$, whose columns $\hat{\bm{\beta}}_{k}$ are defined by (4), satisfies
    $\left\|P(\widehat{\mathbf{B}})-P({\mathbf{B}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\sqrt{\dfrac{L^{*}}{\lambda_{d}}}+L^{*}\right).$
@@
    If Condition 4.7 also holds and at least one tuning parameter $\mu_{2k}$ satisfies
    $\min_{j\in\mathcal{S}^{k}}|\beta_{jk}|>\frac{2\xi_{k}\hat{\rho}_{k}}{2-\alpha}\varphi\mu_{2k},$
    where $\xi_{k}=\|\Sigma_{\widehat{\mathbf{X}}}\bm{\beta}_{k}\|_{2}$ and $\hat{\rho}_{k}={\bm{\eta}}_{k}^{\top}\widehat{\bm{\eta}}_{k}$, $k=1,\ldots,d$, then $\widehat{\mathcal{S}}=\mathcal{S}$ with the same probability,.
Parts (a) and (b) in Theorem 4 for multiple index models essentially require Theorems 2 and 3, respectively, to hold for every dimension. As such, the upper bound on the estimation error of $P(\widehat{\mathbf{B}})$ depends on the lowest non-zero eigenvalue $\lambda_{d}$ of $\bm{\Lambda}$, and requires all sample eigenvectors $\widehat{\bm{\eta}}_{k}$ to not be orthogonal to their corresponding eigenvector $\bm{\eta}_{k}$ for $k=1,\ldots,d$. Under these conditions though, we can establish bounds for the estimation and selection consistency for the endogeneous multiple index model (2).

### v2

@@ 4.1 First stage results
In the first stage of the two-stage lasso SIR procedure, each column of the matrix $\bm{\Gamma}$ is a lasso estimator from the regression of $\mathbf{x}_{j}$ on the design matrix $\mathcal{Z}$. For completeness, we re-state the following result from Lin et al., (2015), which itself is a direct application of Banerjee et al., (2014)
@@
Assume Conditions 3 and 4 are satisfied, and set
@@
Then with probability at least $1-\exp(-C^{\prime}\log q)$, the columns of $\widehat{\bm{\Gamma}}$ satisfy $\lVert\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\rVert_{1}\leq C\sigma_{\max}r\{\log(pq)/n\}^{1/2},$ and with probability at least $1-\exp(-C^{\prime}\log pq)$, we obtain
@@ 4.2 Second stage results
In the second stage, we compute the lasso SIR estimator using the outcome $y$ and the fitted value matrix $\widehat{\mathcal{X}}=\mathcal{Z}\widehat{\bm{\Gamma}}$ obtained from the first stage. Since $\mathbf{X}$ is related to the instruments $\mathbf{Z}$ via (2), then we have $E(\widehat{\mathbf{X}}\mid y)={\bm{\Gamma}}^{\top}E(\mathbf{Z}\mid y)$, where $E(\mathbf{Z}\mid y)$ is related to the dimension reduction subspace $\mathcal{S}_{y\mid\mathbf{Z}}$ of $y$ on $\mathbf{Z}$. It follows that we can impose technical conditions similar to those in Lin et al., (2019) to ensure this subspace can be estimated consistently in high-dimensional settings:
Condition 6.
There exist constants $C_{\min}$ and $C_{\max}$ such that $0<C_{\min}\leq\lambda_{q}(\bm{\Sigma}_{\mathbf{Z}})\leq\cdots\leq\lambda_{1}(\bm{\Sigma}_{\mathbf{Z}})<C_{\max}<\infty$.
Condition 7.
@@
There exist positive constants $c$ and $\kappa\geq 1$ such that $0<c<\lambda_{d}(\bm{\Omega})\leq\ldots\leq\lambda_{1}(\bm{\Omega})\leq\kappa\lambda_{d}(\bm{\Omega})<\lambda_{1}(\bm{\Sigma}_{\mathbf{Z}})$, where $\bm{\Omega}=\operatorname{\text{cov}}\left\{E\left({\mathbf{Z}}\mid y\right)\right\}$, and $n\lambda_{1}(\bm{\Omega})=q^{\alpha}$ for some $\alpha>1/2$.
Condition 8.
The quantity ${m}_{\mathbf{Z}}(y)=E(\mathbf{Z}\mid y)$ satisfies the sliced stability condition introduced in Lin et al., (2018).
@@
Condition 6 ensures the loss function corresponding to the second stage of the proposed lasso SIR estimator does not have too small curvature, and is similar to a condition imposed by Lin et al., (2015) for the two-stage penalized least square method. Condition 7 is analogous to the coverage condition imposed in the SIR literature for $\bm{\Lambda}$ (e.g., Li,, 1991). Finally, Condition 8 controls the smoothness of the conditional expectation ${m}_{\mathbf{Z}}(y)=E(\mathbf{Z}\mid y)$. Such a condition is similarly used in Tan et al., (2018) when no endogeneity is present, with analogous conditions on the total variation having been employed earlier by Hsing and Carroll, (1992); Zhu et al., (2006) for other sufficient dimension reduction methods.
@@
We next establish the consistency of the two-stage lasso SIR estimator for $\text{span}(\mathbf{B})$. First, consider the case of $d=1$ i.e., the single index model version of (2). Here, Lemma 1 implies the covariance matrix $\bm{\Lambda}=\operatorname{\text{cov}}\{E(\widehat{\mathbf{X}}\mid y)\}$ has rank $d=1$ and hence can be written in the form $\bm{\Lambda}=\lambda\bm{\eta}\bm{\eta}^{\top}$, where $\lambda>0$ denotes the non-zero eigenvalue of $\bm{\Lambda}$ and $\bm{\eta}$ is the corresponding eigenvector. Without loss of generality, assume $\|\bm{\beta}\|=1$. For convenience below, we omit the subscript $k$ and use $\hat{\lambda}$ and $\hat{\bm{\eta}}$ to denote the largest eigenvalue and its corresponding eigenvector of $\hat{\bm{\Lambda}}$, respectively, noting both were defined above (4). Furthermore, we write $\tilde{\mathbf{y}}_{1}=\tilde{\mathbf{y}}$. With this notation, the two-stage lasso SIR estimator for the column space of $\bm{\beta}$ is given by
@@
By Lemma (1), we have $\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}=\xi\bm{\eta}$, where $\xi=\lVert\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}\rVert_{2}>0$ is a constant. Letting $\tilde{\bm{\eta}}=P(\bm{\eta})\hat{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\hat{\bm{\eta}}=\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}$, where $\tilde{\bm{\beta}}=\xi^{-1}\bm{\beta}{\bm{\eta}}^{\top}\hat{\bm{\eta}}$, then it is straightforward to see $\tilde{\bm{\beta}}$ has the same direction as $\bm{\beta}$, i.e., $\mathcal{P}(\tilde{\bm{\beta}})=\mathcal{P}(\bm{\beta})$, when $\bm{\eta}^{\top}\hat{\bm{\eta}}\neq 0$, which occurs with probability one. With this quantity, we obtain the following deterministic result.
@@
Suppose the sample covariance matrix $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}=n^{-1}\widehat{\mathcal{X}}^{\top}\widehat{\mathcal{X}}$ satisfies an restricted eigenvalue property with parameters $(\kappa_{2},3)$; see Proposition 11 in Appendix B that shows the latter holds with high probability. If the tuning parameter satisfies $\mu_{2}\geq 2\left\|\hat{\bm{\eta}}-n^{-1}\hat{\mathcal{X}}^{\top}\hat{\mathcal{X}}\tilde{\bm{\beta}}\right\|_{\infty}$, then any solution $\hat{\bm{\beta}}$ defined in (8) satisfies
$\lVert\hat{\bm{\beta}}-\tilde{\bm{\beta}}\rVert_{2}\leq\dfrac{12}{\kappa_{2}}\mu_{2}\sqrt{s}.$
Following Lemma 2, by selecting a value of $\mu_{2}$ so that the condition $\mu_{2}\geq 2\left\|\hat{\bm{\eta}}-n^{-1}\hat{\mathcal{X}}^{\top}\hat{\mathcal{X}}\tilde{\bm{\beta}}\right\|_{\infty}$ is satisfied with high probability, we obtain the consistency of the two-stage lasso SIR estimator for the single index model.
@@
Assume Conditions 3-8 are satisfied, and choose the first stage tuning parameters $\mu_{1j}$ as in (6). Let $\bm{\eta}$ be the eigenvector associated with the largest eigenvalue of $\bm{\Omega}=\text{Cov}\{E(\mathbf{Z}\mid y)\}$. Let $\rho_{1}=L^{2}+\lambda^{-1/2}{L(1+\Psi_{1})}$ and $\rho_{2}=L\sigma_{\max}+\lambda^{-1/2}(1+\Psi_{1})\Psi_{2}$, where $\Psi_{1}=\left\|\mathbf{P}\right\|_{1}$ and $\mathbf{P}$ is the projection matrix associated with the space spanned by $\bm{\eta}\bm{\Gamma}$ and $\Psi_{2}=\|\bm{\eta}\bm{\eta}^{\top}\|_{1}$. If we select the second stage tuning parameters as
$\mu_{2}=C\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right),$ (9)
then with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$, the two-stage lasso SIR estimator $\hat{\bm{\beta}}$ satisfies
$\left\|P(\hat{\bm{\beta}})-P({\bm{\beta}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right).$
Compared with the bound for the one-stage lasso SIR estimator, which is established by Lin et al., (2019), the theoretical bound for the two-stage lasso SIR estimator has an additional factor of order $\left\{\log(pq)/n\right\}^{1/2}$, reflecting the consequences of endogeneity. Also, compared with the result for the two-stage penalized least squares estimator in the linear outcome model of Lin et al., (2015), the bound in Theorem 2 involves an additional term $\rho_{1}\left(\log q/n\right)^{1/2}$, which reflects the consequences of non-linearity. Intuitively, the estimation error of $\hat{\bm{\beta}}$ depends on the estimation error of $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$, which in turn depends upon the estimation error of both the coefficients $\hat{\bm{\Gamma}}$ in the first stage and the covariance matrix $\bm{\Sigma}_{\mathbf{Z}}$.
Next, we establish the selection consistency of the two-stage lasso SIR estimator. Let $\hat{S}=\{j:\hat{\beta}_{j}\neq 0\}$, and partition the matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}=\bm{\Gamma}^{\top}\bm{\Sigma}_{\mathbf{Z}}\bm{\Gamma}$ as
@@
Let $\phi=\|(\mathbf{C}_{SS})^{-1}\|_{\infty}$, and note in this case we have $\xi=\lVert\bm{\Sigma}_{\widehat{\mathbf{X}}}\bm{\beta}\rVert_{2}=\lVert\mathbf{C}_{SS}\bm{\beta}_{S}\rVert_{2}$. We require an additional condition as follows.
Condition 9.
@@
Condition 9 is similar to the mutual incoherence condition for the ordinary lasso estimator in the linear model without endogeneity (e.g., see Section 7.5.1, Wainwright,, 2019).
@@
Assume Conditions 3-9 are satisfied, and choose the first stage tuning parameters $\mu_{1j}$ as in (6). If there exists a constant $C$ such that
@@
and the second stage tuning parameter is selected to satisfy
$\min_{j\in S}|\beta_{j}|>\frac{2\xi\bm{\eta}^{\top}\widehat{\bm{\eta}}}{2-\alpha}\varphi\mu_{2},$ (11)
then with probability at least $1-\exp(C^{\prime}\log q)-\exp(C^{\prime\prime}\log pq)$, the support of the two-stage lasso SIR estimator $\hat{\bm{\beta}}$ satisfies $\hat{S}=S$.
Since the mutual incoherence condition 9 is imposed for the population quantity $\bm{\Sigma}_{\widehat{\mathbf{X}}}$, the additional assumption in (10) is necessary to ensure that this condition is satisfied by the sample covariance matrix $\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}$ with high probability. Also, (11) has an additional factor $\xi\bm{\eta}^{\top}\widehat{\bm{\eta}}$ in the numerator, reflecting another consequence of non-linearity. On the other hand, the lower bound is small when $\alpha$ is small i.e., the lower the correlation between the important and non-important variables, the smaller the magnitude of non-zero components which can be identified.
@@ 4.3 Extension to multiple index models
We generalize the above results to multiple index models. For $k=1,\ldots,d$, let $\widehat{{S}}_{k}=\{j:|\hat{\beta}_{jk}|\neq 0\}$ denote the estimated support set for the vector $\hat{\bm{\beta}}_{k}$, and $\widehat{{S}}=\cup_{k}{\mathcal{S}_{k}}$. Recall $\mathbf{B}$ is only identifiable up to rotational transformation, and the latter only preserves the zero rows of $\mathbf{B}$. As such, selection in the multiple index model refers only to selecting non-zero rows of $\mathbf{B}$. Equivalently, the set $\hat{{S}}$ should be a consistent estimator of the support set $S$ for the diagonal elements of $\mathcal{P}(\mathbf{B})$.
@@
Assume Conditions 3-8 are satisfied, and choose the first stage tuning parameters $\mu_{1j}$ as in (6). Let $\rho_{1}=L^{2}+\lambda_{d}^{-1/2}L(1+\Psi_{1})$ and $\rho_{2}=L\sigma_{\max}+\lambda_{d}^{-1/2}(1+\Psi_{1})\Psi_{2}$, where $\lambda_{d}$ is the $d$th largest eigenvalue of $\Lambda=\operatorname{\text{cov}}\{E(\widehat{\mathbf{X}}\mid y)\}$. If we select the second stage tuning parameters to satisfy
$\mu_{2k}=C\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right),\;k=1,\ldots,d,$ (12)
@@
    Then with probability at least $1-d\exp(-C^{\prime}\log q)-d\exp(-C^{\prime\prime}\log pq)$, the two-stage lasso SIR estimator $\widehat{\mathbf{B}}=(\hat{\bm{\beta}}_{1},\ldots,\hat{\bm{\beta}}_{d})$, whose columns $\hat{\bm{\beta}}_{k}$ are defined by (4), satisfies
    $\left\|P(\widehat{\mathbf{B}})-P({\mathbf{B}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right).$
@@
    If Condition 9 also holds and at least one tuning parameter $\mu_{2k}$ satisfies
    $\min_{j\in\mathcal{S}^{k}}|\beta_{jk}|>\frac{2\xi_{k}{\bm{\eta}}_{k}^{\top}\widehat{\bm{\eta}}_{k}}{2-\alpha}\varphi\mu_{2k},$
    where $\xi_{k}=\|\Sigma_{\widehat{\mathbf{X}}}\bm{\beta}_{k}\|_{2}$, then $\widehat{\mathcal{S}}=\mathcal{S}$ with the same probability,.
[…]